# The late Miocene 10Be anomaly and the possibility of a supernova

### v1

@@ Acknowledgements.
EM, JA, and AR were co-funded by the European Union (ERC, ISM-FLOW, 101055318). JG was co-funded from the European Union, Central Bohemian Region, and Czech Academy of Sciences, as part of the MERIT fellowship (MSCA-COFUND Horizon Europe, Grant agreement 101081195). This work has made use of data from the European Space Agency (ESA) mission Gaia https://www.cosmos.esa.int/gaia, processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement.References
@@
- Cantat-Gaudin et al. (2020) Cantat-Gaudin, T., Anders, F., Castro-Ginard, A., et al. 2020, A&A, 640, A1
@@
- O’Neill et al. (2024) O’Neill, T. J., Zucker, C., Goodman, A. A., & Edenhofer, G. 2024, arXiv e-prints, arXiv:2403.04961
@@
- Sim et al. (2019) Sim, G., Lee, S. H., Ann, H. B., & Kim, S. 2019, Journal of Korean Astronomical Society, 52, 145
@@
- Wallner et al. (2015a) Wallner, A., Bichler, M., Buczak, K., et al. 2015a, Phys. Rev. Lett., 114, 041101
- Wallner et al. (2015b) Wallner, A., Faestermann, T., Feige, J., et al. 2015b, Nature Communications, 6, 5956
@@
- Zucker et al. (2025) Zucker, C., Redfield, S., Starecheski, S., Konietzka, R., & Linsky, J. L. 2025, arXiv e-prints, arXiv:2504.00093
@@ Appendix A Cluster catalog data
In this section, we describe the compilation of open clusters used in this study and explain how we refined the positional and kinematic data obtained from the primary catalog. As outlined in Sect. 2, we primary used the open cluster catalog by Hunt & Reffert (2023), which currently represents the largest open cluster search conducted in a homogeneous way. This catalog was constructed using the Hierarchical Density-Based Spatial Clustering of Applications with Noise routine (HDBSCAN; McInnes et al. 2017) on Gaia DR3 astrometric data (Gaia Collaboration et al. 2023), combined with a statistical density test and a Bayesian convolutional neural network for result validation (see Hunt & Reffert 2023, for details). The catalog contains a total of 7166 star clusters. Following Maconi et al. (2025), we added four additional clusters: CWNU 1028, NGC 1977, OC 0340, and UBC 207. These clusters were identified within other cluster catalogs (Sim et al. 2019; Liu & Pang 2019; Cantat-Gaudin et al. 2020; Szilágyi et al. 2021; Hao et al. 2022; He et al. 2022) which predated Hunt & Reffert (2023) and are considered to be part of the Radcliffe Wave (Konietzka et al. 2024; Maconi et al. 2025).
@@
For six clusters in the Orion region (ASCC 19, ASCC 20, OCSN 56, OCSN 61, OCSN 65, and Theia 13), we were able to update the stellar membership using the SigMA clustering algorithm (Significance Mode Analysis; Ratzenböck et al. 2023a), applied to Gaia data and specifically tuned for the Orion complex (A. Rottensteiner, in preparation). The updated stellar membership for these six Orion clusters, compared to Hunt & Reffert (2023), changes as follows: ASCC 19 from 61 to 796, ASCC 20 from 194 to 525, OCSN 61 from 147 to 530, OCSN 65 from 70 to 534, OCSN 56 from 88 to 52, and Theia 13 from 249 to 221. The latter two have slightly fewer stellar members compared to Hunt & Reffert (2023). This can be explained by the fact that Hunt & Reffert (2023) prioritized precision over completeness, aiming to minimize the number of false positives associated with each cluster and to build a uniform catalog across the entire Solar neighborhood. In contrast, the updated stellar memberships were derived by applying SigMA specifically to the Orion region. We refer to Ratzenböck et al. (2023a), for a discussion on systematic biases coming from various clustering methods applied to the same region. The initial cluster sample comprises a total of 7170 members.
@@
To obtain more accurate 3D space motions for each cluster, we supplement the Gaia DR3 RVs with additional RV data from supplementary surveys, including APOGEE-2 DR17 (Abdurro’uf et al. 2022), GALAH DR4 (Buder et al. 2025), RAVE DR6 (Steinmetz et al. 2020), Gaia ESO DR6 (Randich et al. 2022), two RV compilations (Gontcharov 2006; Torres et al. 2006), and LAMOST DR5 (Zhao et al. 2012, as corrected by Tsantaki et al. 2022). In cases where a star has multiple RV measurements from different surveys, we selected the RV value with the lowest uncertainty. Additionally, we only include sources with RV errors smaller than . To remove outliers, we applied 3-sigma-clipping around each clusters median RV value. We then derived the Cartesian velocities relative to the Sun () using the sub-sample of stars with valid RV measurements. For our study, we used medians in and , along with the associated uncertainties, as an estimate for the clusters bulk position and motion. Only clusters with at least three stars with RV measurements are included in our final sample, reducing the initial sample from 7170 to 2725.
@@ Appendix B Cluster properties
In this Section we report the Tables that summarize the properties of the clusters as derived in the main part of the paper. The tables regards the 19 clusters that have a probability greater than 1% of hosting a SN within 100 pc of the Solar System and during the time interval 11.5–10.1 Myr ago, as described in Sects. 3 and 4.
@@ Appendix C Initial condition test
We tested the robustness of our result by varying the initial solar parameters (i.e. the Sun’s height above the disk, Galactocentric distance, and velocity respect to the Local Standard of Rest; LSR), as there is no unique definition for these quantities. We then integrated the orbits of the Solar System and the clusters and performed the same analysis described in Sect. 3.
For this test, we assumed a Galactocentric distance of (GRAVITY Collaboration et al. 2018), a vertical height of (Bennett & Bovy 2019), and a solar velocity relative to the LSR of () = (11.1, 12.24, 7.25) (Schönrich et al. 2010). We find that the past relative distances of the Sun and the clusters, and thus the resulting probabilities of having a SN within the 10Be anomaly window, remain largely unchanged over the past 20 Myr. This can be explained by the relatively short integration time and by the fact that we consider the relative distances between the clusters and the Sun rather than their absolute positions. This is also in agreement with previous studies that used similar integration periods (see e.g., Miret-Roig et al. 2020), further supporting the robustness of our conclusions.

### v2

@@ Acknowledgements.
We thank the referee A. Wallner for the insightful comments, which have improved the quality and readability of the paper. EM, JA, and AR were co-funded by the European Union (ERC, ISM-FLOW, 101055318). JG was co-funded by the European Union, Central Bohemian Region, and Czech Academy of Sciences, as part of the MERIT fellowship (MSCA-COFUND Horizon Europe, Grant agreement 101081195). This work has made use of data from the European Space Agency (ESA) mission Gaia https://www.cosmos.esa.int/gaia, processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement.References
@@
- Bally (2008) Bally, J. 2008, in Handbook of Star Forming Regions, Volume I, ed. B. Reipurth, Vol. 4, 459
@@
- Korschinek et al. (2010) Korschinek, G., Bergmaier, A., Faestermann, T., et al. 2010, Nuclear Instruments and Methods in Physics Research B, 268, 187
@@
- Kutschera (2013) Kutschera, W. 2013, International Journal of Mass Spectrometry, 349-350, 203
@@
- O’Neill et al. (2024) O’Neill, T. J., Zucker, C., Goodman, A. A., & Edenhofer, G. 2024, ApJ, 973, 136
@@
- Rugel et al. (2009) Rugel, G., Faestermann, T., Knie, K., et al. 2009, Phys. Rev. Lett., 103, 072502
@@
- Schulreich et al. (2023) Schulreich, M. M., Feige, J., & Breitschwerdt, D. 2023, A&A, 680, A39
@@
- Wallner et al. (2015) Wallner, A., Bichler, M., Buczak, K., et al. 2015, Phys. Rev. Lett., 114, 041101
- Wallner et al. (2016) Wallner, A., Feige, J., Kinoshita, N., et al. 2016, Nature, 532, 69
@@
- Zucker et al. (2025) Zucker, C., Redfield, S., Starecheski, S., Konietzka, R., & Linsky, J. L. 2025, ApJ, 986, 58
@@ Appendix A Cluster catalog data
In this section, we describe the compilation of open clusters used in this study and explain how we refined the positional and kinematic data obtained from the primary catalog. As outlined in Sect. 2, we primarily used the open cluster catalog by Hunt & Reffert (2023), which currently represents the largest open cluster search conducted homogeneously. This catalog was constructed using the Hierarchical Density-Based Spatial Clustering of Applications with Noise routine (HDBSCAN; McInnes et al. 2017) on Gaia DR3 astrometric data (Gaia Collaboration et al. 2023), combined with a statistical density test and a Bayesian convolutional neural network for result validation (see Hunt & Reffert 2023, for details). The catalog contains a total of 7166 star clusters. Following Maconi et al. (2025), we added four additional clusters that are associated with the Radcliffe wave (see also Konietzka et al. 2024): CWNU 1028, NGC 1977, OC 0340, and UBC 207.
@@
For six clusters in the Orion region (ASCC 19, ASCC 20, OCSN 56, OCSN 61, OCSN 65, and Theia 13), we were able to update the stellar membership using the SigMA clustering algorithm (Significance Mode Analysis; Ratzenböck et al. 2023a), applied to Gaia data and specifically tuned for the Orion complex (A. Rottensteiner, in preparation). The updated stellar membership for these six Orion clusters, compared to Hunt & Reffert (2023), changes as follows: ASCC 19 from 61 to 796, ASCC 20 from 194 to 525, OCSN 61 from 147 to 530, OCSN 65 from 70 to 534, OCSN 56 from 88 to 52, and Theia 13 from 249 to 221. The latter two have slightly fewer stellar members compared to Hunt & Reffert (2023). The fact that the SigMA-selected clusters are generally richer than those identified by Hunt & Reffert (2023) can be explained by the different selection strategies. The latter study, which focuses on a larger stellar population sample, applies more stringent criteria and prioritizes precision over completeness to minimize the number of false positives associated with each cluster and to construct a uniform catalog across the entire Solar neighborhood. In contrast, the updated stellar memberships presented here are derived by applying SigMA specifically to the Orion region, allowing for a more detailed analysis. We refer to Ratzenböck et al. (2023a) for a discussion on systematic biases coming from various clustering methods applied to the same region. The initial cluster sample comprises a total of 7170 members. In Appendix E, we evaluate the impact of the membership lists on the SN probability estimation.
@@
To obtain more accurate 3D space motions for each cluster, we supplement the Gaia DR3 RVs with additional RV data from supplementary surveys, including APOGEE-2 DR17 (Abdurro’uf et al. 2022), GALAH DR4 (Buder et al. 2025), RAVE DR6 (Steinmetz et al. 2020), Gaia ESO DR6 (Randich et al. 2022), two RV compilations (Gontcharov 2006; Torres et al. 2006), and LAMOST DR10 or LAMOST DR5 (Zhao et al. 2012; Cui et al. 2012) (DR5 as corrected by Tsantaki et al. 2022 from the so called SoS catalog. SoS-LAMOST-DR5 was only used if the source was not in LAMOST DR10). In cases where a star has multiple RV measurements from different surveys, we selected the RV value with the lowest uncertainty. Additionally, we only include sources with RV errors smaller than . To remove outliers, we applied 3-sigma clipping around each cluster’s median RV value. We then derived the Heliocentric Galactic Cartesian velocities () [km/s] using the sub-sample of stars with valid RV measurements. For our study, we used medians in and , along with the associated uncertainties, as an estimate for the cluster’s bulk position and motion. Only clusters with at least three stars with RV measurements are included in our final sample, reducing the initial sample from 7170 to 2725.
@@ Appendix B Cluster properties
In this Section, we report the Tables that summarize the properties of the clusters as derived in the main part of the paper. The tables regard the 19 clusters that have a probability greater than 1% of hosting an SN within 100 pc of the Solar System and during the time interval 11.5–10.1 Myr ago, as described in Sects. 3 and 4.
@@ Appendix C Initial condition test
We tested the robustness of our result by varying the initial solar parameters (i.e., the Sun’s height above the disk, Galactocentric distance, and velocity with respect to the LSR), as there is no unique definition for these quantities. We then integrated the orbits of the Solar System and the clusters and performed the same analysis described in Sect. 3.
For this test, we assumed a Galactocentric distance of (GRAVITY Collaboration et al. 2018), a vertical height of (Bennett & Bovy 2019), and a solar velocity relative to the LSR of () = (11.1, 12.24, 7.25) (Schönrich et al. 2010). We find that the past relative distances of the Sun and the clusters, and thus the resulting probabilities of having an SN within the 10Be anomaly window, remain largely unchanged over the past 20 Myr. This can be explained by the relatively short integration time and by the fact that we consider the relative distances between the clusters and the Sun rather than their absolute positions. This is also in agreement with previous studies that used similar integration periods (see e.g., Miret-Roig et al. 2020), further supporting the robustness of our conclusions.
@@ Appendix D Systematic error study
Appendix E Probability using the original cluster stellar memberships
In this Section, we present the results for the SN probability computed using the original stellar memberships for the Orion clusters. As described in Sect. 3 and in the Appendix A, our main analysis is based on the catalog by Hunt & Reffert (2023), to which we:
added the four additional clusters CWNU 1028, NGC 1977, OC 0340, and UBC 207;
updated the stellar membership list of the Orion clusters ASCC 19 (from 61 to 796), ASCC 20 (from 194 to 525), OCSN 56 (88 to 52), OCSN 61 (147 to 530), OCSN 65 (70 to 534), and Theia 13 (249 to 221). Among these clusters, some of them (e.g., ASCC 20, OCSN 56, OCSN 61) have a significant contribution to the total SN probability.
Here, to test the impact of the membership lists, we recomputed the probability of a close SN within the 10Be anomaly using the original cluster’s memberships as provided by Hunt & Reffert (2023). Following the same procedure described in Sect. 3, we determined the cluster ages, applied corrections for mass incompleteness, and estimated both the time and distance to the Solar System of the SNe. The results of this analysis are presented in Fig. 3 and in Table 4.
We found that:
among the Orion’s clusters, ASCC 20, OCSN 56, and OCSN 61 still represent the main contributors to the SN probability;
for ASCC 20 and OCSN 61, the probabilities are lower than those reported in the main analysis due to the smaller number of original stellar members in these clusters. Nevertheless, they remain the main contributors to the total SN probability;
the contribution of ASCC 20 to the SN probability decreases from 23.0% to 12.0% at 100 pc. However, it still dominates at distances below 50 pc, with a probability of 8.9% (compared to 13.5% in the main analysis);
the contribution of OCSN 61 shows the largest decrease, from 29% to 9.4% at 100 pc. With the original membership list, however, its contribution begins at a closer distance (about 50 pc instead of 70 pc). This difference arises because the cluster’s velocity and position change slightly when considering the original members rather than the updated ones;
the contribution of OCSN 56 increases, since its original number of stellar members is greater than in the main analysis. Its probability rises from 3.7% to 7.3% at 100 pc;
the total SN probability as estimated from all the clusters decreases from 68.0% to 55.4% at 100 pc, and from 13.9% to 10.1% at 50 pc. For more details, see Table 4.
From this analysis, we conclude that the probability of an SN during the 10Be anomaly is lower than that found in our main analysis, but still high enough to consider this scenario possible.
| All clusters* | 55.4 | 48.3 | 41.1 | 34.6 | 29.0 | 24.6 | 21.5 | 18.8 | 16.2 | 13.1 | 10.1 | 7.7 | 5.4 | 2.7 | 0.5 |
| ASCC 19 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | - | - | - |
| ASCC 19* | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | - | - |
| ASCC 20* | 12.0 | 12.0 | 12.0 | 12.0 | 12.0 | 12.0 | 12.0 | 11.9 | 11.5 | 10.4 | 8.9 | 7.3 | 5.3 | 2.7 | 0.5 |
| OCSN 56* | 7.3 | 5.9 | 4.9 | 3.2 | 2.3 | 1.1 | 0.6 | 0.2 | 0.1 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| OCSN 61* | 9.4 | 9.4 | 9.3 | 9.1 | 8.7 | 8.0 | 6.9 | 5.5 | 4.0 | 2.4 | 1.2 | 0.3 | 0.1 | 0.0 | 0.0 |
| OCSN 65 | 0.9 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | - | - | - | - | - | - | - |
| OCSN 65* | 1.6 | 0.8 | 0.7 | 0.4 | 0.1 | 0.1 | 0.0 | 0.0 | - | - | - | - | - | - | - |
| Theia 13 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| Theia 13* | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |